Complete the table using the given values of x and the equation y = 4x - 3. Show your work.
x y
-1
1
3

Answer by alicealc(293)

y = 4x - 3

x = -1
y = 4x - 3 = 4*(-1) - 3 = -4 - 3 = -7

x = 1
y = 4x - 3 = 4*1 - 3 = 4 - 3 = 1

x = 3
y = 4x - 3 = 4*3 - 3 = 12 - 3 = 9

So, the complete table is:
x, y
-1, -7
1, 1
3, 9